# Physics-Informed Wireless Imaging with Implicit Neural Representation in RIS-Aided ISAC System

### v1

@@ III INR-Based Wireless Imaging
This section first presents the INR-based imaging algorithm design, and then compares it with traditional model-based and data-driven imaging methods.
@@ III-A CSI Measurements and Forward Model
INR is developed under the supervision of CSI measurements and forward models. This subsection explains how both components are obtained in the considered ISAC system.
@@ III-A1 CSI Measurements
In this study, ${\mathbf{H}}_{\text{sen},k}$ is first estimated before performing imaging. Using the least squares (LS) channel estimation algorithm, $N_{\text{t}}$ received signals are required to derive the estimated CSI, which is given as
@@
where $\mathbf{N}_{\text{sen},k}$ denotes the channel estimation error. As a result, $N_{\text{s}}=KN_{\text{t}}$ symbol intervals are used to acquire sufficient CSI measurements $\{\widehat{\mathbf{H}}_{\text{sen},k}\}_{k=1}^{K}$ for target imaging, corresponding to the $K$ varying RIS phase configurations. By stacking all elements in $\{\widehat{\mathbf{H}}_{\text{sen},k}\}_{k=1}^{K}$, we derive the measurement vector $\mathbf{y}\in\mathbb{C}^{N_{\text{m}}\times 1}$ with the dimension of $N_{\text{m}}=KN_{\text{t}}N_{\text{r}}$. In contrast to prior studies [5, 13], this approach eliminates the need for multipath extraction and simplifies system design.
Figure 2: Illustration of the INR network structure and the relationship between the image grid and the ROI image.
@@ III-A2 Forward Model
The physical relationship between CSI measurements and the ROI image must be leveraged to impose constraints during image reconstruction. We assume that all point-to-point channels remain constant under the quasi-static condition. In (2), ${\mathbf{H}}_{\text{others},k}$ is treated as disturbance to a partially known physical model [13, 5]. Accordingly, (5) can be reformulated as
@@
Consequently, the forward model corresponding to the $k$-th RIS configuration $\bm{\omega}_{k}$ is expressed as $f_{k}(\bm{\sigma},\bm{\omega}_{k})=\text{vec}({\mathbf{H}}^{\prime}_{\text{sen},k})$, where $\text{vec}(\cdot)$ denotes the vectorization of a matrix. By stacking all $K$ components $\{f_{k}(\bm{\sigma},\bm{\omega}_{k})\}_{k=1}^{K}$, the complete forward model is denoted as $f(\bm{\sigma},\bm{\Omega})$.
@@ III-B Network Structure
INR employs an NN to learn the representation of the continuous shape and scattering characteristics of the ROI image. Following advanced DL methods [10], the input to the INR is any position $\mathbf{p}=[x,y,z]^{\text{T}}$ within the ROI, and its output is the corresponding image parameter at that location, i.e., the scattering coefficient at $\mathbf{p}$, denoted as $\sigma_{\mathbf{p}}\in\mathbb{R}$. Let the INR be denoted by $\mathcal{M}_{\bm{\theta}}$ with trainable parameters $\bm{\theta}$; then, we have
@@
Thus, the continuous ROI image is implicitly stored in $\mathcal{M}_{\bm{\theta}}$ through its specific NN architecture and parameters.
We adopt a six-layer multi-layer perceptron (MLP) for the INR, with 256 neurons per layer. However, MLPs tend to exhibit a bias toward learning low-frequency functions, making it difficult to capture high-frequency details presented in ROI images, such as sharp target edges where small movements in position $\mathbf{p}$ cause significant voxel value changes. To overcome this limitation, we introduce specific positional encoding and a designated activation function to improve the MLP’s ability to represent these high-frequency components.
@@ III-B1 Positional Encoding
The 3D position $\mathbf{p}$ is not a suitable direct input to the MLP, since $\mathbf{p}$ is low-dimensional and lacks high-frequency components beneficial for INR-based imaging. Thus, $\mathbf{p}$ should be transformed into a higher-dimensional representation that captures more frequency components. This process is referred to as positional encoding. In this study, we employ Fourier feature mapping, defined as [14]
@@
where Fourier features are represented by $\sin$ and $\cos$ functions. $\mathbf{B}\in\mathbb{R}^{N_{\text{pe}}\times 3}$ is a random Gaussian matrix, with elements independently sampled from a Gaussian distribution $\mathcal{N}(0,\chi^{2})$. The randomness in $\mathbf{B}$ enables $\bm{\gamma}(\mathbf{p})$ to include components across a broad frequency spectrum. Here, $N_{\text{pe}}$ is the dimensionality of positional encoding, and $\chi$ is a hyperparameter controlling the variance of the elements in $\mathbf{B}$.
@@ III-B2 Activation Function
The commonly used rectified linear unit (ReLU) performs poorly in learning high-frequency signals, due to its piecewise linear nature and lack of second-order derivatives. Alternatively, continuous and periodic functions have been shown to be effective as activation functions in INR-based architectures [15]. In this study, we replace ReLU with the $\sin$ function to improve the MLP’s ability to learn high-frequency content. The $i$-th hidden layer can be expressed as
@@
where $\mathbf{x}_{i}$, $\mathbf{W}_{i}$, and $\mathbf{b}_{i}$ represent the input, weights, and biases of the $i$-th layer, respectively. $\kappa$ is a hyperparameter controlling the periodicity of the sine activation. As a result, the continuous image features and high-frequency components are better preserved and recovered within the MLP. Moreover, we apply the NN parameter initialization strategy proposed in [15] to tackle the high sensitivity of initialization caused by sine-based non-linearities. Finally, the output layer employs the Sigmoid activation function, and the overall INR network architecture is illustrated in Fig. 2.
@@ III-C NN Training and Image Generation
By storing image information in the NN, INR transforms direct image-domain optimization into parameter-domain optimization, given by
$\text{(P1)}\quad\bm{\theta}^{\star}=\mathop{\arg\!\min}\limits_{\bm{\theta}}E(f(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega}),\mathbf{y})+\rho(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ})),\\$
where $\bm{\mathcal{G}}^{\circ}\in\mathbb{R}^{N_{\text{v}}\times 3}$ denotes the discrete sampling grid of the continuous image. $\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ})$ represents the predicted ROI image of the INR, whose corresponding CSI is calculated by $f(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega})$. $E(f(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega}),\mathbf{y})$ measures the error between $f(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega})$ and CSI measurements $\mathbf{y}$, and $\rho(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}))$ is the regularization term capturing the image prior. Moreover, Problem (P1) also leverages an implicit image-statistics prior encoded by the NN structure $\mathcal{M}_{\bm{\theta}^{\star}}$, meaning that all position-parameter pairs $(\mathbf{p},\sigma_{\mathbf{p}})$ share a common continuous underlying function [10].
The training procedure iteratively follows four steps:
@@ III-C1 Step 1—NN Inference
In each epoch, the NN renders the image stored in its parameters. A sampling grid $\bm{\mathcal{G}}\in\mathbb{R}^{N_{\text{g}}\times 3}$ is generated in advance, containing $N_{\text{g}}$ point locations within the ROI. This grid may be generated through uniform or non-uniform sampling in the 3D ROI space. Sampling density can be adapted using priors to form denser grids in regions of interest or sparser grids in empty areas [10]. [^1] Since INR learns a continuous representation of the image, the final output resolution $\bm{\mathcal{G}}^{\circ}$ can be determined without the limitation of the training grid $\bm{\mathcal{G}}$. The NN receives the grid locations $\bm{\mathcal{G}}$ as input and outputs the corresponding predicted scattering coefficients:
@@
By rearranging the elements in $\widehat{\bm{\sigma}}_{\bm{\theta}}$ based on their spatial positions, an ROI image is generated, whose resolution is determined by the sampling density of $\bm{\mathcal{G}}$.
@@ III-C3 Step 3—Loss Calculation
The NN is trained to minimize the discrepancy between the predicted CSI $\widehat{\mathbf{y}}_{\bm{\theta}}$ and the measured CSI $\mathbf{y}$. The physics-informed loss function is defined as
@@
where the first term enforces CSI consistency (the data term in (P1)), and the second term is an explicit regularizer promoting sparsity in the ROI image. The hyperparameter $\alpha$ controls the weight of the sparsity constraint.
@@ III-C4 Step 4—NN Parameter Update
where $r$ is the learning rate and $\nabla_{\bm{\theta}}L_{\bm{\theta}}$ denotes the gradient with respect to $\bm{\theta}$. This update can be implemented using standard DL frameworks such as PyTorch or TensorFlow.
In summary, INR performs imaging using only the forward model and CSI measurements, similar to model-based approaches such as FT and CS. However, the optimization of the ROI image is realized through DL techniques, akin to data-driven methods. After the NN is trained, the final ROI image can be rendered by reapplying Step 1 with the desired image grid $\bm{\mathcal{G}}^{\circ}$ as input, given as
$\widehat{\bm{\sigma}}=\mathcal{M}_{\bm{\theta}^{\star}}(\bm{\mathcal{G}}^{\circ}).$ (15)
Although the training grid $\bm{\mathcal{G}}$ is discretized, the NN implicitly learns and stores a continuous representation of the ROI image. Therefore, INR allows flexible output image resolution by adjusting the sampling density of $\bm{\mathcal{G}}^{\circ}$, which is infeasible for traditional model-based or data-based imaging approaches.

### v2

@@ III INR-Based Wireless Imaging
INR is a novel technique grounded in physical modeling and implemented using DL techniques. This section first presents the essential CSI measurements and forward model used by INR. Then, the network structure and INR-based imaging algorithm design are presented subsequently. Finally, the advantages of INR are highlighted by comparing it with traditional model-based and data-driven DL methods.
@@ III-A CSI Measurements and Forward Model
INR is developed under the supervision of CSI measurements and forward models, as described in (1). This subsection explains how they are obtained in the considered ISAC system.
@@ III-A1 CSI Measurements
In this study, ${\mathbf{H}}_{\text{sen},k}$ is first estimated before performing imaging. Under the least squares (LS) channel estimation algorithm, $N_{\text{t}}$ signals are required [45].  [^1] The number of measurements required to estimate ${\mathbf{H}}_{\text{sen},k}$ can be reduced well below $N_{\text{t}}$ by leveraging the sparse nature of ${\mathbf{H}}_{\text{sen},k}$ [45]. In this study, we adopt the LS method as a simple example. The transmit signals are denoted as $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N_{\text{t}}}]$, which are known at the BS and can be designed using dedicated precoding techniques [7]. The estimated CSI is then expressed as
@@
where $\mathbf{N}_{\text{sen},k}$ denotes additional noise caused by channel estimation errors. As a result, $N_{0}=KN_{\text{t}}$ symbol intervals are used to acquire sufficient CSI measurements $\{\widehat{\mathbf{H}}_{\text{sen},k}\}_{k=1}^{K}$. Based on the general model in (1), the measurement vector $\mathbf{y}$ is formed by stacking all elements in $\{\widehat{\mathbf{H}}_{\text{sen},k}\}_{k=1}^{K}$ into a single vector of dimension $N_{\text{m}}=KN_{\text{t}}N_{\text{r}}$. In contrast to prior studies [18, 17, 9], this approach eliminates the need for multipath extraction and simplifies system design.
@@ III-A2 Forward Model
The physical relationship between CSI measurements and the ROI image must be leveraged to impose constraints during image reconstruction. We assume that all point-to-point channels remain constant under the quasi-static condition. In (10), the term ${\mathbf{H}}_{\text{others},k}$ represents unknown channel components due to limited knowledge of the environment, whereas other components are explicitly modeled in (11) and (12). Thus, ${\mathbf{H}}_{\text{others},k}$ is treated as disturbance to a partially known physical model [17, 18, 9]. Accordingly, (13) can be reformulated as
@@
Consequently, the forward model corresponding to the $k$-th RIS configuration $\bm{\omega}_{k}$ is expressed as $f_{k}(\bm{\sigma},\bm{\omega}_{k})=\text{vec}({\mathbf{H}}^{\prime}_{\text{sen},k})$, where $\text{vec}(\cdot)$ denotes the vectorization of a matrix. By stacking all $K$ components in $\{f_{k}(\bm{\sigma},\bm{\omega}_{k})\}_{k=1}^{K}$, the complete forward model is denoted as $f(\bm{\sigma},\bm{\Omega})$. Note that the forward physical model can be modified with available environmental information, e.g., occlusions and reflections.
Remark 1.
The proposed INR-based wireless imaging method is potentially generalizable to diverse imaging systems, provided that their corresponding CSI measurements and forward physical models are available. The only requirement for physical models is differentiable. To illustrate this applicability, this study investigates the wireless imaging problem within the context of RIS-aided ISAC systems.
@@ III-B Network Structure
The INR technique employs an NN to learn the representation of the continuous shape and scattering characteristics of the ROI image. Following advanced DL methods [24], the input to the INR is any position $\mathbf{p}=[x,y,z]^{\text{T}}$ within the ROI, and its output is the corresponding image parameter at that location, i.e., the scattering coefficient at $\mathbf{p}$, denoted as $\sigma_{\mathbf{p}}\in\mathbb{R}$. Let the INR be denoted by $\mathcal{M}_{\bm{\theta}}$ with trainable parameters $\bm{\theta}$; then, we have
@@
Thus, the continuous ROI image is implicitly embedded into $\mathcal{M}_{\bm{\theta}}$ through its specific NN architecture and parameters.
The network structure of the INR is typically a multi-layer perceptron (MLP). In this study, we adopt an MLP consisting of six fully connected (FC) layers, each containing 256 neurons. However, MLPs tend to exhibit a bias toward learning low-frequency functions, making it difficult to capture high-frequency variations [46]. Note that the term high-/low-frequency here refers to the variation rate of voxel values with respect to their locations, rather than the frequency of wireless signals. The ROI image typically includes a large number of high-frequency components, especially near targets’ edges, where small movements in position $\mathbf{p}$ can lead to significant changes in voxel values. To address this limitation, specific positional encoding and a designated activation function are introduced to enhance the MLP’s capability in learning the continuous ROI image [46].
Figure 4: Illustration of the INR network structure and the online training process.
@@ III-B1 Positional Encoding
Although the INR maps the 3D position $\mathbf{p}$ to its corresponding image parameter, $\mathbf{p}$ is not a suitable direct input to the MLP for two reasons. First, MLPs are generally more effective at learning mappings from high-dimensional inputs to low-dimensional outputs, while $\mathbf{p}$ is low-dimensional and thus contains limited information. Second, position variation in the image is inherently low-frequency, meaning $\mathbf{p}$ itself lacks the high-frequency components beneficial for INR-based imaging. Therefore, $\mathbf{p}$ should be transformed into a higher-dimensional representation that captures more frequency components. This process is referred to as positional encoding.
Although various positional encoding techniques have been utilized in literature [47], this work does not aim to provide a comprehensive review. Instead, we adopt an effective method tailored to the wireless imaging scenario. Specifically, we employ Fourier feature mapping [48], defined as
@@
where Fourier features are represented by $\sin$ and $\cos$ functions. The matrix $\mathbf{B}\in\mathbb{R}^{N_{\text{pe}}\times 3}$ is a random Gaussian matrix, with elements independently sampled from a Gaussian distribution $\mathcal{N}(0,\chi^{2})$. The randomness in $\mathbf{B}$ enables $\bm{\gamma}(\mathbf{p})$ to include components across a broad frequency spectrum. Here, $N_{\text{pe}}$ is the dimensionality of positional encoding, and $\chi$ is a hyperparameter controlling the variance of the elements in $\mathbf{B}$.
@@ III-B2 Activation Function
The rectified linear unit (ReLU) is commonly used as the activation function for FC layers. However, due to its piecewise linear nature and lack of second-order derivatives, ReLU performs poorly in learning high-frequency signals. For instance, it is not well suited to approximating a sine function. Alternatively, continuous and periodic functions have been shown to be effective as activation functions in INR-based architectures [49]. In this study, we replace ReLU with the $\sin$ function to improve the MLP’s ability to learn high-frequency content. The $i$-th hidden layer can be expressed as
@@
where $\mathbf{x}_{i}$, $\mathbf{W}_{i}$, and $\mathbf{b}_{i}$ represent the input, weights, and biases of the $i$-th layer, respectively. The parameter $\kappa$ is a hyperparameter controlling the periodicity of the sine activation. As a result, the continuous image features and high-frequency components are better preserved and recovered within the MLP. However, sine-based non-linearities can lead to high sensitivity to the NN’s parameter initialization. Therefore, we apply the initialization strategy proposed in [49]. Finally, the output layer employs the Sigmoid activation function, and the overall INR network architecture is illustrated in Fig. 4.
@@ III-C NN Training and Image Generation
INR extracts image information from CSI measurements with the aid of the physical model through NN training, where this information is implicitly embedded into NN parameters, rather than directly or explicitly presented as an image with discrete voxels. To achieve this, INR transforms the direct image-domain optimization in Problem (P1) into an indirect parameter-domain optimization, given by
$\text{(P2)}\quad\bm{\theta}^{\star}=\mathop{\arg\!\min}\limits_{\bm{\theta}}\mu(f(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}),\bm{\Omega}),\mathbf{y})+\rho(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ})),\\$
where $\bm{\mathcal{G}}^{\circ}\in\mathbb{R}^{N_{\text{v}}\times 3}$ denotes the discrete sampling grid of the ROI and determines the resolution of the final predicted image.
Based on (P2), the NN is trained to represent the ROI image under the constraints imposed by the CSI measurements $\mathbf{y}$ and the forward model $f(\bm{\sigma},\bm{\Omega})$. In addition to the explicit regularization term $\rho(\mathcal{M}_{\bm{\theta}}(\bm{\mathcal{G}}^{\circ}))$, Problem (P2) also leverages an implicit image-statistics prior encoded by the NN structure $\mathcal{M}_{\bm{\theta}}$, meaning that all position-parameter pairs $(\mathbf{p},\sigma_{\mathbf{p}})$ share a common continuous underlying function [35, 24]. Diverse from traditional data-driven DL methods, the NN training and inference procedures of the proposed INR-based imaging algorithm are both executed in an online manner. Specifically, the online training procedure iteratively follows four steps:
@@ III-C1 Step 1—NN Inference
In each epoch, INR transforms the image information implicitly embedded in the NN parameters to an explicit image by performing inference on the training grid $\bm{\mathcal{G}}\in\mathbb{R}^{N_{\text{g}}\times 3}$, which contains $N_{\text{g}}$ point locations within the ROI. [^2] In practical implementation, we can set $\bm{\mathcal{G}}=\bm{\mathcal{G}}^{\circ}$ and $N_{\text{g}}=N_{\text{v}}$ to use the same image resolution for NN training and final image prediction. This uniform grid can be calculated based on the known ROI boundaries and selected sampling interval [9, 16, 17]. However, $\bm{\mathcal{G}}$ can be different from $\bm{\mathcal{G}}^{\circ}$, since INR employs a continuous function, i.e., the MLP, to represent the ROI image, thus allowing arbitrary sampling grid density. A dense grid $\bm{\mathcal{G}}$ enables the NN to learned fine-grained image details, whereas a sparse grid can accelerate NN training. Moreover, $\bm{\mathcal{G}}$ can be non-uniformly sampled to form denser grids in regions of targets or sparser grids in empty areas [24]. As illustrated by the dashed boxes in Fig. 4, the NN receives the grid locations $\bm{\mathcal{G}}$ as the input and outputs the corresponding predicted scattering coefficients:
@@
By rearranging the elements in $\widehat{\bm{\sigma}}_{\bm{\theta}}$ based on their spatial positions defined by $\bm{\mathcal{G}}$, an ROI image is generated, whose resolution equals the sampling interval of $\bm{\mathcal{G}}$. Note that the NN can generate an image even without training, since its input $\bm{\mathcal{G}}$ has been determined. However, the derived image may be significantly distorted and represents little GT image information, due to under-optimized NN parameters. During training, the NN repetitively generates predicted images in each epoch, which is indirectly supervised by CSI measurements in subsequent steps.
@@ III-C2 Step 2—CSI Prediction
$\widehat{\mathbf{y}}_{\bm{\theta}}$ is the CSI if $\widehat{\bm{\sigma}}_{\bm{\theta}}$ was the GT image. This embeds physics domain knowledge into the learning procedure [22, 29].
@@ III-C3 Step 3—Loss Calculation
The NN is trained to minimize the discrepancy between the CSI $\widehat{\mathbf{y}}_{\bm{\theta}}$ corresponding to the predicted image $\widehat{\bm{\sigma}}_{\bm{\theta}}$ and the CSI $\mathbf{y}$ measured by the ISAC system. The loss function is defined as
@@
where the first term enforces CSI consistency (the data term in (P2)), and the second term is an explicit regularizer promoting sparsity in the ROI image. The hyperparameter $\alpha$ controls the weight of the sparsity constraint. Here, the NN output in (19) is not directly supervised, since its GT value is unknown; instead, it is transformed from the image domain to the CSI domain in (20) and indirectly supervised by CSI measurements in (21).
TABLE I: Comparison between model-based FT, CS, data-driven DL, and INR imaging techniques.
FT [11, 18] CS [9, 17, 16, 8, 10] Data-driven DL [19, 20, 21] INR
Measurement amount Large, satisfying Nyquist criterion Small, using sparse prior Small, using training data prior Small, using sparse prior and DL-based optimization
Multipath extraction Request Request None None
Forward model requirement Spatial integral Sensing matrix None Any differentiable model
Image resolution dependence FFT points Predefined voxel gird in the sensing matrix NN structure and training data Arbitrary
“CSI-image” dataset amount None None Extremely Large None
Imaging target requirement None Sparse Same distribution with training data None
Computational platform CPU CPU GPU GPU
Imaging procedure FFT and Hadamard product Iterative CS algorithm Offline training and online inference Online training and image generation
@@ III-C4 Step 4—NN Parameter Update
where $r$ is the learning rate and $\nabla_{\bm{\theta}}L_{\bm{\theta}}$ denotes the gradient of $L_{\bm{\theta}}$ with respect to $\bm{\theta}$. This update can be implemented using standard DL frameworks such as PyTorch or TensorFlow.
The above online training steps have been depicted in Fig. 4, requesting only one piece of CSI measurements $\mathbf{y}$. Moreover, Fig. 4 has been simplified into the dashed box of Fig. 5(c), together with the following online image generation procedure.
After the NN is trained, its parameters have been optimized to implicitly represent the ROI image. To transform this implicitly embedded information into explicit images, we can reapply Step 1 in Sec. III-C1 using the desired image grid $\bm{\mathcal{G}}^{\circ}$ as input. This image generation procedure is given as [^3] The CSI measurements $\mathbf{y}$ and the forward physical model $f(\bm{\sigma},\bm{\Omega})$ are used to assist in NN training. After the training procedure, their associated image information has been extracted and embedded into NN parameters. Thus, the final image generation step does not utilize $\mathbf{y}$ and $f(\bm{\sigma},\bm{\Omega})$.
$\widehat{\bm{\sigma}}=\mathcal{M}_{\bm{\theta}^{\star}}(\bm{\mathcal{G}}^{\circ}),$ (23)
where $\bm{\theta}^{\star}$ denote optimized NN parameters. Although the training grid $\bm{\mathcal{G}}$ is discretized, the NN implicitly learns and encodes a continuous representation of the ROI image. Therefore, INR allows flexible output image resolutions by adjusting the sampling density of $\bm{\mathcal{G}}^{\circ}$ during final image generation. Thus, voxel values off the training grid $\bm{\mathcal{G}}$ can be obtained, which is infeasible for traditional model-based or data-driven imaging approaches. This ability is presented by the simulation results in Sec. VI-B7.